please show the steps to solve this: 11-p=19 thanks!
Found 2 solutions by prince tayem, amoresroy:
Answer by prince tayem(3)
Answer by amoresroy(361)

please show the steps to solve this: 11-p=19 thanks!
subtract 11 to both sides of the equation
11-11-p = 19-11
0-p = 8
-p = 8
divide both sides by -1 to get p
-p/-1 = 8/-1
p = -8